Find the polynomial equation over R of lowest degree which is satisfied by 1-i and 3+2i
1
Expert's answer
2019-07-05T12:53:00-0400
"(x-1+i)(x-1-i)(x-3-2i)(x-3+2i)=0"

"(x^2-2x+2)(x^2-6x+13)=0"

"x^4-8x^3+27x^2-38x+26=0"